OpenStudy (anonymous):

-9(-y – 13) = -13y + 29

OpenStudy (mathlegend):

You need to distribute here because the negative 9 is touching those parentheses

OpenStudy (mathlegend):

So do that first @ur-only_angel

OpenStudy (anonymous):

@MathLegend, what if it wasn't touching the parantheses? What if the -9 and the parantheses were spaced from each other?

OpenStudy (anonymous):

9y + 117 = -13y + 29 Then you solve for y by putting the values with y on one side and without a variable on the other: 9y + 117 -117 = -13y + 29 -117 9y = -13y - 88 9y + 13y = -13y + 13y - 88 22y = -88 22y/22 = -88/22 y = -4 I think...

OpenStudy (mathlegend):

@genius12 if it is not touching then there should be a sign in between

OpenStudy (anonymous):

@MathLegend I was joking, I think u missed it =.=

OpenStudy (mathlegend):

I guess so. :D

OpenStudy (mathlegend):

@LoveYou*69 I do understand that you worked out the problem. However, I would prefer if we would instead let the OP do the work. This way we can check for understanding.